If the mean of 5, 7, x ,10 ,5 and 7 is 7, then x = __________.

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
Mean=(a1+a2+a3+a4+a5)/5 =7

Substituting the given values of a1 to a5,

Mean

7=(5+7+x+10+5)/5

35=27+x

x=8.